# Hard Contacts with Soft Gradients: Refining Differentiable Simulators for Learning and Control

### v1

@@ Abstract
Contact forces pose a major challenge for gradient-based optimization of robot dynamics as they introduce jumps in the system’s velocities. Penalty-based simulators, such as MuJoCo, simplify gradient computation by softening the contact forces. However, realistically simulating hard contacts requires very stiff contact settings, which leads to incorrect gradients when using automatic differentiation. On the other hand, using non-stiff settings strongly increases the sim-to-real gap. We analyze the contact computation of penalty-based simulators to identify the causes of gradient errors. Then, we propose DiffMJX, which combines adaptive integration with MuJoCo XLA, to notably improve gradient quality in the presence of hard contacts. Finally, we address a key limitation of contact gradients: they vanish when objects do not touch. To overcome this, we introduce Contacts From Distance (CFD), a mechanism that enables the simulator to generate informative contact gradients even before objects are in contact. To preserve physical realism, we apply CFD only in the backward pass using a straight-through-trick, allowing us to compute useful gradients without modifying the forward simulation.
Project page: https://sites.google.com/view/diffmjx
Keywords: Differentiable Simulation, Contact Modeling, Motion Planning, MPC, System Identification, Dexterous Manipulation.
@@ 1 Introduction
$**{}^{\textbf{*}}$ Equal contribution.
Simulators have proven to be an indispensable tool for synthesizing policies for robot control. Sim-to-real techniques are now the de facto standard in creating robust and performant policies for real robots [1, 2, 3, 4, 5]. Predominantly, these simulators have been used as black-box functions where policies are optimized with reinforcement learning (RL). If simulators provided informative gradients for actions and simulation parameters, a set of very powerful tools and applications would become available. We could generate policies with gradient-based MPC and RL for high-dimensional control problems that are much more sample-efficient than with classical deep RL. We could optimize the simulator to fit real-world observations to bridge the sim-to-real gap via gradient-based system identification. With both in place, policies for new tasks could be generated and deployed in seconds instead of hours.
What keeps us from doing it? Current simulators already provide gradients, but they are either too slow, incorrect, or uninformative. We care about sufficiently realistic simulations for real robot applications, which implies hard materials and stiff contact models. The principal source of problems is that simulators only compute a discrete approximation $\widetilde{F}$ of the continuous physics $F$. As a result, the gradient $\nabla\widetilde{F}$ (computed using automatic differentiation) may not accurately represent the true gradient $\nabla F$. In the sequel, gradient “correctness” refers to how closely this gradient matches the gradient of the continuous model. We show that in penalty-based simulators, which are the standard [6] in multibody physics simulation, the gradients for dynamics with contacts are incorrect to the extent that their sign can be flipped. In addition, the simulations need to be fast, so we cannot resort to infinitesimally small simulation timesteps. Previous suggestions for time-of-impact correction [7, 8] in elastic simulators do not solve the problem for penalty-based collision models, whereas softening contact dynamics or resorting to unfeasibly small timesteps do improve gradients. We propose using adaptive timestep integrators under the hood while keeping the fixed-timestep external interface, creating a small computational overhead but yielding correct gradients, even for hard contact cases.
@@
Another obstacle for gradient-based optimization for policy generation and system identification is the non-informativeness of gradients about unmade contacts. For example, when a robot’s hand is not in contact with an object, then there is no gradient directing it to make contact for task facilitation. Therefore, drawing inspiration from prior works on contact-invariant optimization [9, 10], we propose Contacts From Distance (CFD) to address this problem. However, when done naively, introducing artificial contact forces considerably changes the simulation, resulting in a too large sim-to-real gap. In order to preserve the initial simulation, we propose using the straight-through-trick to only introduce CFD in the gradient computation (backward pass). We implement our changes in Mujoco XLA, the JAX [11] implementation of MuJoCo [12], and additionally fix some low-level collision routines to be truly differentiable. As a result, our implementation makes use of GPU acceleration and automatic differentiation. We show in proof-of-concept simulations that we can perform system identification and policy synthesis for collision-rich problems and hard contact in high-dimensional systems such as musculoskeletal models shown in Fig. 1.
Figure 2: Left: Common computational graph for robot control synthesis. Center: If the initial ball radius $p$ is set too small, then the gradient $\nabla_{\!p}L$ resulting from a one-step-ahead prediction is zero. Right: Large constraint penetration for hard contacts results in large gradients $\nabla_{\!x_{k}}L$ which aggravates the estimation of the previous state $x_{k}$.
2 Computing gradients in penalty-based simulators
@@
As illustrated in Fig. 2, we want to use automatic differentiation to obtain the correct gradient of a loss functional $L(\tilde{x}_{k+1},x_{k},a_{k},p)$ where the next state of the robotic system is governed by the discrete-time dynamics $x_{k+1}=\mathrm{step}(x_{k},a_{k},p)$ with the state $x_{k}:=x(t_{k})=[q_{k},v_{k}]$ at time $t_{k}$ consisting of the system’s generalized position $q_{k}\in\mathbb{R}^{n_{q}}$ and velocity $v_{k}\in\mathbb{R}^{n_{v}}$, control actions $a_{k}\in\mathbb{R}^{n_{a}}$, and model parameters $p\in\mathbb{R}^{n_{p}}$. Typically, multi-body dynamics simulators consist of forward dynamics model and a numerical integration method. The forward dynamics govern the system’s acceleration $\dot{v}$ via the equations of motion
@@
with the joint-space inertia matrix $M(q)\in\mathbb{R}^{n_{v}\times n_{v}}$, the applied forces $\tau(x,a)\in\mathbb{R}^{n_{v}}$, the bias force $c(x)\in\mathbb{R}^{n_{v}}$, the constraint space Jacobian $J(q)\in\mathbb{R}^{n_{c}\times n_{v}}$, and the constraint forces $f(x)=f_{\mathcal{E}}+f_{\mathcal{F}}+f_{\mathcal{C}}\in\mathbb{R}^{n_{c}}$ consisting of the equality constraint, the generalized friction, and the contact constraint forces. The Jacobian $J$ maps joint velocity $v$ into constraint space, whereas $J^{\top}$ maps $f$ into generalized coordinate space. The contact-free dynamics are typically derived via recursive multi-body algorithms [13, 14] while computing contact forces requires an intricate interplay of collision detection and contact force computation. Simulators typically use numerical integrators such as semi-implicit Euler and 4th-order Runge-Kutta. As we will see throughout this work, numerical integration plays a critical role for understanding how contact forces may hinder correct gradient computation. For that, we will focus on MuJoCo XLA as a concrete penalty-based simulator.
2.1 MuJoCo XLA
MuJoCo XLA (MJX) is a reimplementation of MuJoCo using the Python library JAX, which enables GPU-parallelizable gradient computation via automatic differentiation. This capability allows MJX to compute gradients of rigid-body dynamics efficiently. MuJoCo has become the de facto standard in robotics, alongside other widely used simulators such as Bullet [15], Drake [16], DiffTaichi [7] and IsaacGym [17]. The growing importance of MuJoCo within the robotics community is further evidenced by a collaborative effort between NVIDIA and Google to develop the general-purpose simulator “Newton”, built on top of the recently introduced MuJoCo Warp [18]. In order to understand where gradients computed in MJX may fail to approximate the correct ones, it is essential to first develop a solid understanding of how MuJoCo computes constraint forces.
@@ 2.1 MuJoCo XLA
A collision detector is an algorithm that, given the state $x$ and geometry parameterizations of two bodies, returns the signed distance $r(x)$ between potential contact point candidates alongside with corresponding body surface normals. Collision detection is critical for contact force computation, as contacts are only considered active – that is, a contact point exerts contact forces – if $r(x)<0$. For a robot simulator to be fully differentiable, it is key that the collision detector is itself differentiable. While collision detection is not the primary focus of this work, we note that obtaining reliable gradients from MJX required addressing several discontinuities in its collision detection module.
@@
When objects in a simulation come into contact, we need to determine not only where and when forces act, but also their direction and strength. MuJoCo solves this using an optimization approach inspired by Gauss’s principle [19, 20], minimizing the deviation from the unconstrained acceleration (what the system would do subject to constraint-free forces alone), while ensuring that all constraints are satisfied. Rather than solving rigid equations exactly, MuJoCo uses a softened formulation that blends two ideas: minimizing unexpected accelerations, and steering the system towards satisfying constraints using spring-damper dynamics.
The core of this approach lies in two functions: the impedance $d(r)$, and the position-level reference acceleration $h(r)$. The impedance $d(r)\in(0,1)$ describes how strongly a constraint resists violation – it controls both how much force can be applied and how costly that force is to the solver. As illustrated in Fig. 3, the impedance $d\in(0,1)$ is a polynomial spline function of the constraint violation $r$ specified by the parameters solimp $=(d_{o},d_{w},w,\text{midpoint},\text{power})$. The reference acceleration $h(r)$, in turn, defines how the solver aims to correct constraint violations over time, based on the velocity and position error, and given solref parameters consisting of the time constant $t_{c}$ and the damping ratio $\phi_{d}$. Internally, MuJoCo reformulates this as a convex optimization problem and solves it efficiently using a Newton method. For an in-depth description of Mujoco’s contact computation consult Suppl. B.1.
@@ 3 Correcting the contact gradients of penalty-based simulation
3.1 Analyzing gradient quality
To evaluate the correctness of gradients in the presence of contacts, we unroll the trajectory of several primitives bouncing against a plane and observe the final position and its gradient with respect to the initial velocity. In Fig. 5 (MJX row), we observe that the gradient is oscillating with an amplitude orders of magnitude larger than the loss.
Point mass example.
To better understand this issue, we simplify the setup even further. We consider a point mass colliding with a flat surface in the absence of gravity, simulated via a minimal version of MuJoCo’s penalty-based contact model, as shown in Fig. 5 (middle column). The loss is the distance between the point’s final state and a target. Depending on the initial start height, it shows a cyclic pattern as before, which causes rapid sign flips in the gradient. Similar oscillations have been discussed in previous works [7, 8]. DiffTaichi [7] describes the “time-of-impact” (TOI) problem, showing an oscillating pattern in gradients during the ideal elastic collision. We replicate this for comparison in our toy simulator Fig. 5 (right column). These oscillations arise from time-discretization errors in ODE integration that also affect penalty-based simulators, but have to be addressed differently.
Figure 4: Toy simulation of a point mass colliding with a surface that either resorts to an ideal-elastic contact model (similar to DiffTaichi) or a penalty-based contact model (similar to MuJoCo). Loss and gradients with respect to $q_{0}$. The loss $L=|q_{N}-1|$ uses the final state $q_{N}$, which is obtained after doing $N$ numerical integration steps starting from $q_{0}$ with $v_{0}=-1$. In the penalty-based simulation, reducing $h$ fixes the gradient oscillations, whereas for the ideal-elastic collision, reducing $h$ does not yield correct gradients.
Figure 5: Simulation of geometric primitives thrown onto a surface. Contacts with stiff contact settings cause MJX’s gradients of the toss distance to deviate from central difference gradients, while DiffMJX maintains close agreement.
@@ 3.1 Analyzing gradient quality
For an ideal elastic collision, the ODE is piecewise linear, and at the contact, the velocity is inverted (see also Figure S4 in the Appendix). The TOI approach [7] dynamically splits the ODE into two linear segments at the time of contact, thereby eliminating the discretization error and yielding correct gradients. In penalty-based simulation, the ODE is also linear before and after the collision, but is non-linear and with variable stiffness over the time of the collision. Therefore, it cannot be easily divided into large linear segments. A possible solution would be to reduce the stepsize until the discretization error decreases enough. This is also confirmed in our toy simulator in Fig. 5: Reducing the stepsize in the penalty-based simulation eventually yields correct gradients, whereas for ideal elastic collision, the gradients remain incorrect.
Unfortunately, simply reducing the step size is not a practical solution, as it necessitates extremely small steps that substantially increase the computational and memory demands of gradient computation. This trade-off raises a critical question: Can we retain realistic contacts while maintaining practical simulation speeds?

### v2

# Differentiable Simulation of Hard Contacts with Soft Gradients for Learning and Control

@@ Abstract
Contact forces introduce discontinuities into robot dynamics that severely limit the use of simulators for gradient-based optimization. Penalty-based simulators such as MuJoCo, soften contact resolution to enable gradient computation. However, realistically simulating hard contacts requires stiff solver settings, which leads to incorrect simulator gradients when using automatic differentiation. Contrarily, using non-stiff settings strongly increases the sim-to-real gap. We analyze penalty-based simulators to pinpoint why gradients degrade under hard contacts. Building on these insights, we propose DiffMJX, which couples adaptive time integration with penalty-based simulation to substantially improve gradient accuracy. A second challenge is that contact gradients vanish when bodies separate. To address this, we introduce contacts from distance (CFD) which combines penalty-based simulation with straight-through estimation. By applying CFD exclusively in the backward pass, we obtain informative pre-contact gradients while retaining physical realism.
Project page: https://github.com/martius-lab/diffmjx
@@ 1 Introduction
Gradients have powered major advances in machine learning ranging from video language models to robot control. In robotics, imitation and reinforcement learning widely rely on gradient-based optimization. Yet, despite the dominance of sim-to-real techniques that leverage robot simulators for policy learning  (Tan et al., 2018; Lee et al., 2020; and, 2020; Radosavovic et al., 2024; Li et al., 2022), most methods conspicuously avoid using simulator gradients. This is a missed opportunity considering that simulator gradients offer a direct route to updating actions and learning model parameters. If these gradients were accurate, we could fit simulators to real data significantly narrowing the sim-to-real gap and accelerate policy learning – enabling policies for new tasks to be trained in seconds rather than hours. Given the utility of simulator gradients, what prevents their use in robot learning? In practice, two fundamental issues hinder the use of simulator gradients: (i) discontinuities arising from contacts yield erroneous gradients, and (ii), if objects do not touch, contact gradients are zero. In this work, we chart a path that tackles both challenges.
Differentiating through contacts.
The choice of contact model has significant implications on a simulator’s differentiability. Complementary-based solvers, such as Taylor et al. (2022), compute contact forces exactly, such that sudden jumps in the dynamics aggravate gradient computation. Hence, recent literature on differentiating through complementary-based solvers proposes analytical reformulations of the dynamics, e.g. based on the implicit function theorem (Werling et al., 2021; Taylor et al., 2022), or resorts to randomized smoothing (Tassa & Todorov, 2010; Duchi et al., 2012; Suh et al., 2022; Bouyarmane et al., 2009; Xu et al., 2010). While differing in computational cost, both approaches have similar empirical performance (Pang et al., 2023; Schwarke et al., 2024). To improve computational efficiency, MuJoCo (Todorov, 2014) reformulates complementary constraints as a convex optimization problem, where a constraint’s ability to generate force grows proportionally to the constraint violation. While such a penalty-based simulator can be made smooth, gradient inaccuracies increase with the contact stiffness, the relative contact velocity, and the integration step size. While common advice points to reducing the step size to improve gradient accuracy, using sufficiently small step sizes results in prohibitively slow simulation.
Computing gradients of stiff simulations.
In the sequel, we show that gradient errors in penalty-based simulators arise from computing a discrete approximation of continuous physics – put simply, from numerically integrating stiff differential equations. Consequently, the gradients obtained via automatic differentiation are “incorrect” insofar as they do not align with those of the underlying continuous system. Although adaptive integration is well studied, it’s rarely used in robotics and to the best of our knowledge its utility for differentiable simulation has not been explored. This perspective complements prior proposals for time-of-impact correction (Hu et al., 2020; Schwarke et al., 2024) in impulse-based simulators. We incorporate adaptive timestep integration into MuJoCo XLA, incurring a slight computational overhead while obtaining correct gradients in the presence of hard contacts, and remaining compatible with existing MuJoCo libraries.
Figure 1: Common computational graph for robot control synthesis.
Computing gradients between non-colliding objects.
@@
Another obstacle for gradient-based optimization for policy generation and system identification is the non-informativeness of gradients about unrealized contacts. For example, when a robot’s hand is not in contact with an object, then there is no gradient directing it to make contact for task facilitation. Therefore, drawing inspiration from prior works on contact-invariant optimization (Mordatch et al., 2012b; a), we propose Contacts From Distance (CFD) to address this challenge. However, naively introducing artificial contact forces considerably alters the simulation, resulting in a too large sim-to-real gap. In order to preserve the simulation realism, we propose using the straight-through-trick to introduce CFD solely in the gradient computation. We implement our changes in Mujoco XLA, and additionally fix some low-level collision routines to be truly differentiable. CFD complements position-based costs often used in robot learning by removing the need to specify exact contact locations.
In summary, DiffMJX and CFD provide additional tuning knobs to enable practitioners to set gradient correctness and inform an optimization on how non-colliding objects would need to move to influence each other’s state. In turn, this work provides a new perspective on how to obtain useful gradients in penalty-based differentiable simulators thereby enabling parameter estimation and policy synthesis for collision-rich and high-dimensional systems.
2 Robot simulation
@@
As illustrated in Fig.˜1, we want to use automatic differentiation to obtain the correct gradients of a loss functional $L(\tilde{x}_{k+1},x_{k},a_{k},p)$ where the next state of the robotic system is governed by the discrete-time dynamics $x_{k+1}=\mathrm{step}(x_{k},a_{k},p)$ with the state $x_{k}:=x(t_{k})=[q_{k},v_{k}]$ at time $t_{k}$ consisting of the system’s generalized position $q_{k}\in\mathbb{R}^{n_{q}}$ and velocity $v_{k}\in\mathbb{R}^{n_{v}}$, control actions $a_{k}\in\mathbb{R}^{n_{a}}$, and model parameters $p\in\mathbb{R}^{n_{p}}$. Typically, multi-body dynamics simulators consist of forward dynamics model and a numerical integration method. The forward dynamics govern the system’s acceleration $\dot{v}$ via the equations of motion
@@
with the joint-space inertia matrix $M(q)\in\mathbb{R}^{n_{v}\times n_{v}}$, applied forces $\tau(x,a)\in\mathbb{R}^{n_{v}}$, bias force $c(x)\in\mathbb{R}^{n_{v}}$, constraint space Jacobian $J(q)\in\mathbb{R}^{n_{c}\times n_{v}}$, and constraint forces $f(x)=f_{\mathcal{E}}+f_{\mathcal{F}}+f_{\mathcal{C}}\in\mathbb{R}^{n_{c}}$ consisting of the equality constraint, the generalized friction, and the contact constraint forces. The contact-free dynamics are typically derived via recursive multi-body algorithms (Featherstone, 2014) while computing contacts are resolved through an intricate interplay of collision detection and contact force optimization. As we will see throughout this work, numerical integration plays a pivotal role in understanding how contact forces may hinder correct gradient computation. For that, we will resort to MuJoCo XLA as a concrete example of a penalty-based simulator.
2.1 Constraint resolution in MuJoCo XLA
MuJoCo XLA (MJX) is a reimplementation of MuJoCo using the Python library JAX (DeepMind et al., 2020), which enables GPU-parallelizable gradient computation via automatic differentiation. MuJoCo has become the de facto standard in robotics, alongside other widely used simulators such as Bullet (Coumans & Bai, 2016), Drake (Russ Tedrake, 2019), DiffTaichi (Hu et al., 2020) and Nvidia PhysX (Liang et al., 2018). MuJoCo ’s importance is underpinned by NVIDIA and Google recently releasing MuJoCo Warp (Howell, 2025). In what follows, we provide a brief overview of how MuJoCo resolves contacts, which serves as a hands-on example of a penalty-based simulator.
@@ 2.1 Constraint resolution in MuJoCo XLA
Given the state $x$ and geometry parameterizations of two bodies, a collision detector returns the signed distance $r$ between potential contact points alongside with their surface normals. A contact is only considered active – that is, a contact point can exert contact force – if $r<0$. While collision detection is not the primary focus of this work, we later show how to remove discontinuities from its functions to ensure differentiability.
@@
A detailed description of MuJoCo is given in its documentation, here we provide a summary of key equations. Given Eq.˜1, MuJoCo resolves constraints via a relaxation of Gauss’ principle (Gauß, 1829) which in its primal formulation reads
$\begin{split}(\dot{v},\dot{\omega})&=\underset{(x,y)}{\arg\min}\bigl\|x-M^{-1}(\tau-c)\bigr\|_{M}^{2}+\|y-a_{\mathrm{ref}}\|_{R^{-1}}^{\mathrm{H}}\\ \text{subject}\text{ to}&\>J_{\mathcal{E}}x-y_{e}=0,\>J_{\mathcal{F}}x-y_{\mathcal{F}}=0,\>J_{\mathcal{C}}\,x-y_{\mathcal{C}}\in\mathcal{K}^{*}\end{split}$ (2)
with the friction cone (dual) $\mathcal{K}^{*}$, the regularizer $R>0$, and Huber norm $\|\cdot\|^{\mathrm{H}}$. The reference acceleration $a_{\mathrm{ref}}$ denotes the solver’s target for the constraint space acceleration $\dot{\omega}$. Drawing inspiration from Baumgarte (1972), $a_{\mathrm{ref}}$ follows a damped harmonic oscillator
$a_{\mathrm{ref},i}=-b_{i}(Jv)_{i}-k_{i}r_{i}=-\frac{2}{d_{w}\cdot t_{c}}(Jv)_{i}-\frac{d(r_{i})}{d_{w}^{2}\cdot t_{c}^{2}\cdot\phi_{d}^{2}}r_{i}$ (3)
whose dynamics are determined by the impedance $d(r)$, and the solref parameters consisting of the time constant $t_{c}$ and the damping ratio $\phi_{d}$. The impedance $d(r)$ is the central function for determining constraint forces (such as contact forces). As illustrated in Fig.˜2, the impedance $d\in[0,1]$ is a function of the constraint violation $r$ and is specified by the solimp parameters $(d_{o},d_{w},w,\text{midpoint},\text{power})$ which define its shape as a polynomial spline. MuJoCo’s documentation refers to the impedance as a “constraints ability to generate force” as it determines $a_{\mathrm{ref}}$ and weights the cost for applying constraint forces in Eq.˜2 via the diagonal matrix $R$, which is computed as $R_{ii}=\hat{A}_{ii}(1-d_{i})/d_{i}$, where $\hat{A}$ denotes an approximation to $A=\text{diag}(JM^{-1}J^{\top})$. In turn, the constraint is hard if $R\rightarrow 0$, and approaches an infinitely soft (i.e. non-existent) constraint in the limit $R\rightarrow\infty$. If $\mathcal{K}^{*}$ solely contains pyramidal or elliptic cone constraints, then problem (2) reduces to a convex problem that is solved efficiently via an exact Newton method.
@@ 3 Correcting the contact gradients of penalty-based simulation
To evaluate the correctness of gradients in the presence of contacts, as illustrated in the top row of Fig.˜4, we unroll the trajectory of several primitives bouncing against a plane to observe the final position and its gradient with respect to the initial velocity. In Fig.˜4 (MJX row), at the default integration step of 0.002 s, we observe that the loss is oscillating at a high frequency, which results in large fluctuations of the gradient. Similar oscillations have been discussed in previous works (Hu et al., 2020; Schwarke et al., 2024). In particular, Hu et al. (2020) emphasizes that the “time-of-impact” (TOI) causes gradient oscillations when using ideal elastic or complementary collisions. Yet, TOI gradient errors differ in nature from errors observed in penalty-based simulators; as illustrated in the example below. We observe that state oscillations and the corresponding gradient artifacts are not a characteristic of the specific simulator. Instead, more fundamentally, they arise from time-discretization errors in the ODE integration. While this affects both penalty-based and ideal-elastic simulators, we will see that the issue has to be addressed differently.
Example 0 – Point collision:
As illustrated in Fig.˜4 (right), a point mass starts at height $q_{0}$ with velocity $v_{0}=-1$ and is integrated for $N$ steps with semi-implicit Euler in the absence of gravity. It collides with a flat surface and bounces back up. Contact collision is either resolved via a minimal version of a penalty-based contact model as found in MuJoCo or an ideal elastic collision as in (Hu et al., 2020). The corresponding JAX code is shown in Fig.˜16 in the Appendix. The loss $L=|q_{N}-q_{T}|$ is the distance between the point’s final state $q_{N}$ and the target height $q_{T}=1$. For ideal-elastic collisions, we observe sawtooth-like loss oscillations in $L$, resulting in the gradient $\nabla_{q_{0}}L$ with the wrong sign, independently of the stepsize as observed by Hu et al. (2020). For the penalty-based simulation, gradient oscillations notably reduce when lowering the stepsize.
Figure 3: Simulation of primitives thrown onto a surface. For stiff contacts, MJX’s gradients of the toss distance deviate from central differences, while DiffMJX maintains agreement.
Figure 4: Toy simulation of a point mass colliding with a surface that either resorts to an ideal-elastic contact model used in DiffTaichi or a penalty-based contact model similar to MuJoCo. Decreasing the integration stepsize $h$ solely reduces errors in the penalty-based simulation, in the ideal-elastic simulation the gradient sign remains wrong.
@@
For an ideal elastic collision, the ODE of our minimal example is piecewise linear: The dynamics are linear (due to the absence of gravity) before and after the contact, at which the velocity is inverted (see also Fig.˜16 in the Appendix). The TOI approach proposed by Hu et al. (2020) exploits this structure by dynamically splitting the ODE into two linear segments at the time of contact. Integrating those separately thereby eliminates the discretization error and yields correct gradients. In penalty-based simulation, the ODE is linear before and after the collision, but is non-linear with variable stiffness over the time of the collision. Therefore, it cannot be easily divided into large linear segments.
Luckily, for penalty-based simulation we can instead rely on a different technique for reducing the gradient error. The simple solution is to reduce the stepsize, which does not work in the ideal elastic case, as seen in Fig.˜4. From an ODE perspective, this works because the penalty-based ODE is smooth, allowing us to continuously control the integration error by reducing the stepsize. Unfortunately, simply reducing the step size is not a practical solution, as it necessitates extremely small steps that substantially increase the computational and memory demands of gradient computation. This trade-off raises a critical question: Can we retain correct gradients of realistic contacts while maintaining practical simulation speeds?